### v1

# From “Worse is Better” to Better: Lessons from a Mixed Methods Study of Ansible’s Challenges

Abstract.
Infrastructure as Code (IaC) tools have transformed the way IT infrastructure is automated and managed, but their growing adoption has also exposed numerous challenges for practitioners. In this paper, we investigate these challenges through the lens of Ansible, a popular IaC tool. Using a mixed methods approach, we investigate challenges, obstacles, and issues faced by practitioners. We analyze 59,157 posts from Stack Overflow, Reddit, and the Ansible Forum to identify common pain points, complemented by 16 semi-structured interviews with practitioners of varying expertise levels.
Based on our findings, we propose four main recommendations to improve Ansible: 1) refactoring to mitigate performance issues, 2) restructuring higher-level language concepts, 3) improved debugging and error reporting tools, and 4) better documentation and learning resources. By highlighting the real-world struggles of Ansible users, we provide actionable insights for tool designers, educators, and the broader IaC community, contributing to a deeper understanding of the trade-offs inherent in IaC tools.
copyright: none
1. Introduction
Infrastructure as Code (IaC) has emerged as a critical paradigm in modern software engineering, enabling organizations to automate the configuration, provisioning, and management of IT infrastructure. By treating infrastructure configurations as code, IaC enables practitioners to use software engineering practices such as automated testing and version control, and facilitates reproducibility, scalability, and maintainability of infrastructure resources.
IaC has become widespread across industries, with several tools gaining significant adoption. Examples of widely used IaC tools include Ansible (Red Hat, 2024), Chef (Progress, 2024), and Puppet (Perforce, 2024) for configuration management, and Terraform (HashiCorp, 2024) and CloudFormation (Amazon Web Services, Inc., 2011) for infrastructure provisioning. These tools allow organizations to automate complex infrastructure tasks, reducing manual effort and the risk of human error. For example, NASA and Siemens use Ansible to automate infrastructure deployment and configuration management (Red Hat, Inc., 2023; NASA Goddard Space Flight Center, 2023). The widespread adoption of these tools highlights their importance in modern software operations.
Despite the significant advantages of IaC, practitioners face several challenges. Since IaC involves writing code, errors can be introduced, leading to misconfigurations that affect system security (Rahman et al., 2019, 2021; Reis et al., 2022; Saavedra and Ferreira, 2022; Saavedra et al., 2023; Opdebeeck et al., 2023) and system reliability  (GitLab, 2014; Stack Exchange Network, 2014; gooeyblob, 2016; Honeycomb, 2021; Jan David Nose, 2023; Wikimedia, 2017). For example, due to bugs in their IaC scripts, GitHub experienced an outage of their DNS infrastructure (Fryman, 2014) and Amazon Web Services lost around 150 million dollars after issues with their S3 billing system (Hersher, 2017). To address this, there has been an effort by the research community to categorize, identify and repair defects (Saavedra and Ferreira, 2022; Saavedra et al., 2023; Rahman et al., 2021; Reis et al., 2022; Hassan et al., 2024; Begoug et al., 2024; Sotiropoulos et al., 2020; Opdebeeck et al., 2022; Rahman et al., 2023; Lepiller et al., 2021; Schwarz et al., 2018; Sharma et al., 2016; Bessghaier et al., 2024; Opdebeeck et al., 2023; Zerouali et al., 2023; Rahman and Williams, 2018; Rahman et al., 2019, 2020; Saavedra et al., 2025). In addition, IaC tools themselves present challenges, such as unintuitive debugging mechanisms, inconsistent behaviors, and complex learning curves. Prior research has explored challenges faced by IaC practitioners, identifying common issues and limitations of existing tools (Rahman et al., 2018; Guerriero et al., 2019; Begoug et al., 2023; Tanzil et al., 2023). However, existing studies often take a broad perspective, examining challenges across multiple DevOps or IaC tools, which can dilute insights specific to any one tool. Studies focused on a single IaC tool have not, so far, included Ansible.
In this paper, we aim to deepen the understanding of the challenges faced by Ansible users. Ansible is currently the most widely used configuration management tool (Stack Overflow, 2024) and, while it is known that Ansible has issues, such as difficulties with variable scoping (Opdebeeck, 2024) and debugging (Hacker News User gizdan, 2021), the specifics of these challenges and their impact on practitioners remain underexplored. Using a mixed methods approach, we analyze a large corpus of online discussions and conduct semi-structured interviews to systematically examine the obstacles faced by Ansible users. Our study is the first to provide a detailed account of these issues and to offer actionable recommendations for improvement.
In summary, our contributions are:
  • A large-scale empirical analysis of 59,157 online discussions from Stack Overflow, Reddit, and the Ansible Forum to identify common challenges and issues in Ansible.
  • Insights from 16 semi-structured interviews with Ansible practitioners of varying expertise levels, providing qualitative perspectives on usability and technical challenges.
  • Recommendations for improving Ansible, focusing on documentation, debugging tools, and architectural enhancements.
  • Broader implications for IaC tool the design and evolution, highlighting trade-offs between usability, flexibility, and scalability.
Our study aims to inform Ansible’s developers, educators, and the broader IaC community, fostering improvements that enhance the effectiveness and usability of IaC tools. A replication package with all the data and code used in this study is available at:
https://doi.org/10.5281/zenodo.15031088
2. Ansible
According to the Stack Overflow Developers Survey 2024, Ansible is the most widely used configuration management tool (Stack Overflow, 2024).
To define the desired system state, practitioners must write declarative code using Ansible’s hierarchical structure, which consists of four levels: roles, playbooks, plays, and tasks. Roles serve as reusable components that organize related functionality, including tasks, variables, modules, and templates. Playbooks, written in YAML, define configurations by grouping multiple plays. Each play associates specific hosts, identified by tags or hostnames, with a set of tasks. Tasks are the fundamental execution units in Ansible, consisting of an optional name and an action, which are performed using an Ansible module with specified arguments. Modules are reusable, standalone Python scripts that interact with the target host to execute specific tasks. For example, the ansible.builtin.shell module runs shell commands on the host. Ansible also supports blocks that group multiple tasks together.
Execution order. By default, Ansible executes tasks sequentially in the order in which they appear in the playbook. However, certain tasks can be configured to execute only when triggered by other tasks. These tasks, known as handlers, run in response to changes detected in preceding tasks.
Inventory. In Ansible, the hosts to be configured are listed in an inventory, typically stored as an INI or YAML file. In this file, hosts can be grouped by assigning them specific tags. An example of such an inventory is shown in Figure 1.
Variables and Templates. Ansible uses variables to manage system differences, allowing the same variable name to hold different values for each host. These variables can be used in playbooks to adjust task behavior based on the target host. They can be defined in various ways, such as within playbooks or in Ansible’s inventory, as shown in Figure 1. Additionally, Ansible automatically gathers special variables called facts, which provide system information such as operating system details. Variable precedence is determined by the location where the variable is defined and follows a hierarchy of 22 rules (Ansible project contributors, 2025). Ansible leverages Jinja2 templating to enable dynamic expressions and access to variables and facts (Ansible project contributors, 2025), both within the playbooks and in configuration files used via the template module.
Push-based model. Unlike other configuration management tools, such as Chef and Puppet, Ansible follows a push-based configuration model. Typically, configuration management tools use a pull-based model that involves a server that stores configurations and target hosts that retrieve them. However, in a push model, the server actively sends configurations to the target hosts rather than waiting for them to request updates. Since target hosts do not initiate communication, push-based systems such as Ansible do not require an agent on each machine. Instead, a server, commanded by the Ansible user and with SSH access to all hosts, deploys Ansible modules to bring each host to the desired state.
Example. Figure 2 presents excerpts of an Ansible playbook for configuring a web server. This playbook consists of a single play. Here, we show three tasks and one handler. In line 2, the playbook specifies that it applies to all target hosts listed in the inventory. Line 2 defines the use of the webserver role, ensuring that all tasks within this role are executed. Line 2 demonstrates the use of an Ansible fact to display the operating system of the target host, using Jinja2 syntax for variable access, if the condition specified in line 2 is met. In line 2, the playbook specifies that if the state of the nginx package changes, the handler Restart Nginx (line 2) should be triggered. Line 2 employs a Jinja template to dynamically generate a default web page, allowing the use of Ansible variables for customization.
{minted}
[fontsize=]text [all:vars] ansible_user=admin
[web_servers] web1.example.com ansible_host=192.168.1.10 http_port=80 nginx_version=1.18 web2.example.com ansible_host=192.168.1.11 http_port=8080 nginx_version=1.18
Figure 1. Example of an Ansible INI inventory.
{minted}
[linenos=true,fontsize=,xleftmargin=2em, escapeinside=!!]yaml — - name: Set up and configure web server hosts: all !! gather_facts: yes roles: [’webserver’] # Role to set up a web server !! tasks: - name: Display system facts debug: msg: ”The system OS is ansible_facts[’os_family’] .” !! when: ansible_facts[’os_family’] == ’Debian’ !!
- name: Ensure Nginx is installed apt: name: nginx notify: Restart Nginx !!
- name: Deploy an index.html template template: !! src: templates/index.html.j2 dest: /var/www/html/index.html notify: Reload Nginx
(…) handlers: - name: Restart Nginx !! service: name: nginx state: restarted (…)
Figure 2. Excerpts from an Ansible playbook to configure a web server.
3. Mixed Methods Study
In this section, we outline our research questions and describe the methodology used to answer them. Our study follows a mixed methods approach, summarized in Figure 3. First, we collected posts from Q&A platforms and used topic modeling to understand the challenges, obstacles, and issues faced by Ansible practitioners. Then, we conducted semi-structured interviews to gain deeper insights into participants’ experiences with Ansible.
Figure 3. Overview of the research methodology combining automated topic extraction and thematic analysis.
3.1. Research Questions
We aim to answer the following research questions.
  • RQ1: [Issues] What are the common issues practitioners face when using Ansible? In this question, we investigate the common challenges that practitioners face while working with Ansible. By analyzing the emergent topics from practioners’ posts in Q&A platforms and practitioners feedback from interviews, we aim to categorize these challenges and identify patterns that can help us to understand pain points across various praticioner groups.
  • RQ2: [Adoption] What are the aspects that influence the adoption of Ansible? This question focuses on the factors that drive Ansible’s adoption, identified via interviews with practitioners.
  • RQ3: [Improvements] What changes would improve Ansible? This question aims to identify actionable suggestions for improving Ansible. The goal is to provide insights for Ansible’s development community, and broader IaC community, on how to refine the tool to better meet practitioners’ expectations and needs.
3.2. Topic Modeling
Building on previous studies (Begoug et al., 2023; Tanzil et al., 2023), we apply topic modeling techniques to identify the challenges practitioners face when using a specific tool, in our case, Ansible. In particular, we use TopicGPT, a prompt-based framework that uses large language models (LLMs) to uncover latent topics in a text collection (Pham et al., 2024). TopicGPT provides interpretable topics that align better with human categorizations than other approaches such as LDA and BERTopic (Pham et al., 2024). After data is collected, the TopicGPT framework involves three steps: Topic Generation, Topic Refinement, and Topic Assignment. In our study, we also perform manual validation of the Topic Assignment step. We configured TopicGPT to use GPT-4o mini via the OpenAI API, specifically the gpt-4o-mini-2024-07-18 version.
3.2.1. Data Collection
We start by collecting posts about challenges, obstacles, and issues faced by Ansible practitioners from three Q&A platforms: StackOverflow (Stack Exchange, 2025b), Reddit (Reddit, 2025), and the Ansible Forum (Red Hat, 2025a) ( 1 in Figure 3). To collect posts from StackOverflow, we used the Stack Exchange API (Stack Exchange, 2025a) and retrieved all posts with the tag Ansible, which corresponds to a total of 23,000 posts. For Reddit, we used the PullPush API (PullPush team, 2025) to collect all 7,187 posts on the subreddit r/ansible that contained the keyword Ansible in addition to keywords possibly related to issues, such as error, help, or slow. For the Ansible Forum, we used web scraping techniques to collect 28,970 posts in the Get Help and Archives categories. We included the category Archives in our analysis, as it contains all posts from the Ansible Google Group, the predecessor of the Ansible Forum, which were later migrated to the Ansible Forum (Red Hat, 2025b). We filtered out posts related to announcements and releases by the Ansible team. In total, we collected 59,157 posts from the three Q&A platforms, forming our final dataset. Data collection took place between January 24, 2025, and February 1, 2025.
3.2.2. Topic Generation
We use TopicGPT to generate a set of topics given our whole dataset as input ( 2 in Figure 3). A topic consists of a name and a description. For each post, TopicGPT prompts GPT-4o Mini to assign one or more of the already generated topics and generate any new relevant topics that are missing, assigning them accordingly. To increase the probability that the topics generated are relevant to our study, we added to the default prompt of TopicGPT the phrase: “The topics must be related to challenges, issues, or problems that developers face while using Ansible.”. In the end, TopicGPT generated 3,573 topics ( 3 in Figure 3).
3.2.3. Topic Refinement
TopicGPT generates a large number of topics that, while mostly relevant, are too granular for practical analysis. Moreover, Pham et al. (Pham et al., 2024) identified instances of topics that are duplicated or out of scope. TopicGPT performs an automated topic refinement step to overcome these issues. However, in our first experiments, we achieved unsatisfactory results, even when using more powerful LLMs such as o3-mini-2025-01-31. For example, in one of our test runs, this approach was unable to merge the following pair of topics with a cosine similarity of $0.96000.9600$, which TopicGPT uses to measure semantic proximity:
1Regex Issues Desc: Problems related to the use of regular expressions in Ansible tasks.
2Regex Desc: Issues related to the use of regular expressions in Ansible.
For this reason, we adopted an approach similar to TopicGPT, but instead of using LLMs, we relied on manual effort to ensure a more reliable final topic list ( 4 in Figure 3). First, instead of merging the topics first and then removing the infrequent ones, we reversed the order. This approach reduced the number of topic pairs that needed to be analyzed during the merging process. The topic distribution follows a power law, where a small number of topics account for the majority of post attributions, whereas most topics are assigned to only a few posts. To ensure adequate coverage, we initialized an empty set and incrementally added topics in descending order of attribution frequency until reaching 80% of the total attributions. Topics that were not included in the final set were discarded. After this step, our topic list was reduced from 3,573 to 153 topics.
Next, we applied an iterative manual refinement process similar to the merging step used in TopicGPT. In each iteration, we systematically evaluated topic pairs, starting with those exhibiting the highest cosine similarity and proceeding to those with the lowest similarity down to a predefined threshold for that iteration. Similarly to TopicGPT, we used Sentence-Transformer embeddings (Reimers and Gurevych, 2019) to calculate cosine similarity. Then, for each pair, one author determined whether the topics should be merged and, if so, whether they should be combined into one of the existing topics or assigned to a newly created topic. The criterion for merging topics was the presence of significant semantic overlap between them. If a topic was merged in a given iteration, pairs that included the merged topic were excluded from further evaluation within that iteration. A second author independently reviewed the merged topics, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. The process was repeated six times until no further merges were identified, resulting in a candidate topic list. In the first three iterations, we applied a threshold of 0.7, while in the last three iterations, we lowered the threshold to 0.6. After the merging process, we reduced our topic list from 153 to 105 topics.
Finally, following the approaches used in previous studies (Asmussen and Møller, 2019; Jacobi et al., 2018), we manually inspected the topics to identify those relevant to our study. A topic was considered relevant if it is related to the challenges, frustrations, issues, or drawbacks associated with using Ansible. We considered only topics that were specific enough to provide meaningful information about a post. For example, this led us to remove the topic Task Execution with description “Issues related to the execution of tasks in Ansible, including conditions and dependencies for task execution.”. In addition, we refined topic names and descriptions. A topic name was considered improved if it more effectively represented its description, while a description was considered improved if it maintained its core ideas while increasing generality, eliminating unnecessary details, or using a clearer and more concise language. Finally, if we identified any additional pairs of topics that met our merging criteria but were not initially detected during the merging process, we merged them accordingly.
An author reviewed each topic, determining whether it should be removed, renamed, or modified. Subsequently, a second author independently reviewed the proposed changes, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. This process resulted in the final list of topics composed of 87 topics ( 5 in Figure 3).
3.2.4. Topic Assignment
We use the topic list generated in Section 3.2.3 and TopicGPT to assign topics to each post. We also apply the self-correction step provided in the TopicGPT framework (Pham et al., 2024) ( 6 in Figure 3).
3.2.5. Validation
To assess the quality of the topic assignment described in Section 3.2.4, we employ the confirmation labeling approach (Yang et al., 2014). To calculate the precision of topic assignment, this method involves sampling (post, topic) pairs of assignments under evaluation. Rather than presenting raters with the entire taxonomy and asking them to manually assign topics to posts, confirmation labeling simplifies the process by requiring raters to provide binary judgments on whether a given topic is relevant to a particular post. This approach effectively reduces the cognitive overload of a large and complex taxonomy (Yang et al., 2014).
Two authors independently performed confirmation labeling on a sample of 300 (post, topic) pairs ( 7 in Figure 3). The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was approximately 0.29, which according to Landis and Koch’s interpretation (Landis and Koch, 1977), is classified as a fair agreement. A third author performed confirmation labeling on the (post, topic) pairs where the initial two authors’ judgments diverged. The final judgment was decided by majority vote. Ultimately, TopicGPT’s topic assignment achieved a precision ( 8 in Figure 3) of approximately 82.67% (see Supplemental Materials for full topic list).
3.2.6. Analysis
To draw relevant insights across our topic analysis and the interviews, we merged the final list of topics derived from TopicGPT with the thematic codes identified in the interview data (described in Section 3.3). Specifically, we integrated the topics with a set of Challenges codes by aligning each topic with the most closely related theme. This merging process was iterative and followed a double-blind procedure: initially, two coders independently grouped the topics into themes. Finally, both coders met to discuss disagreements and reach a consensus. In analyzing the merged data, we examined the relative frequency of each topic across the resulting themes and investigated how these frequencies evolved. In our analysis, we restricted our scope to the last five years of Q&A data to better capture current and emerging concerns, totaling 23,215 recent posts. In addition to recent posts, we leveraged the data older than five years to identify evolution trends.
3.3. Interviews
In addition to the large-scale analysis of Q&A posts, we conducted semi-structured interviews to gain a deeper understanding of practitioners’ day-to-day experiences with Ansible. Below, we describe the recruitment process, structure of the interviews, and analysis. All interview materials are in the Supplemental Materials.
3.3.1. Participants
We categorize participants based their level of expertise with Ansible. For this we use years of experience due to support in previous work (LaToza and Myers, 2010; Roehm et al., 2012) and also to mitigate the Dunning-Kruger Effect, where users overestimate their skills in self-assessments (Mahmood, 2016). We define three levels of expertise with Ansible: Beginner (less than two years of experience, n = 4); Intermediate (two to five years of experience, inclusive, n = 5); and Experienced (more than five years of experience, n = 7).
Recruitment
We recruited 16 participants through a pre-screening survey distributed via email and various social media platforms, including LinkedIn, X (formerly Twitter), Discord, and Slack ( 1 in Figure 3). To meet our eligibility criteria, participants had to have prior experience using Ansible. Following this stage, participants selected their preferred interview modality: a text-based or video-based session conducted via Zoom. The availability of both formats was intended to address known challenges in recruiting software developers for research studies (Baltes and Diehl, 2016), a challenge we also encountered in our study. A total of 118 out of 120 survey participants were eligible and, among these, 16 were available for an interview (an interview rate of 14%). Most of the participants (11 out of 16) opted for a text interview. Several participants mentioned that they would only be able to participate in the interview if they could do it through text. As an incentive, the interviewed participants could enter a raffle for a \$100 gift card.
The supplementary material contains full details on the demographics of participants. In summary, participants reported anywhere from 1 to more than 9 years of Ansible experience, with many currently using Ansible (12). Most identified as male (15), ranged in age from 18–34 (9), and held at least a bachelor’s degree (13). Almost all had a job in CS and had a formal CS education (15).
Ethical Considerations
Before participating in the survey or the interview, all individuals reviewed the consent form and provided informed consent. This study was reviewed and approved by the ethics board at the first authors’ institution.
3.3.2. Interviews
We conducted semi-structured interviews to obtain a nuanced understanding of the participants’ experiences with Ansible ( 2 in Figure 3). Semi-structured interviews are the most widely used interviewing method (DiCicco-Bloom and Crabtree, 2006) and allow adaptability to explore emerging themes and seek clarifications (Lazar et al., 2017). We piloted the protocol with four participants.Voice-based interviews lasted around 30 minutes. All interviews were recorded and automatically transcribed with participant consent using Zoom’s built-in transcription functionality. Research team members manually reviewed and corrected the resulting transcripts. Each interview followed a structured protocol that encompassed five key thematic areas: contextual background, perceived advantages, challenges, recommendations, and future of Ansible. The interview started with general background questions We then moved the discussion towards the advantages of Ansible and adoption challenges. Then, we focused the discussion on Ansible’s limitations. The interview concluded with exploring recommendations to improve Ansible and future perspectives. We actively sought clarification whenever responses were ambiguous or required elaboration.
3.3.3. Analysis
We developed five emergent codebooks using thematic analysis (Braun and Clarke, 2006; Lazar et al., 2017), following a bottom-up approach that allowed themes, codes to emerge inductively from our data. We started by reading the transcripts to highlight recurring ideas, creating initial codes, and iteratively refining them into coherent themes ( 3 and 4 in Figure 3). Our codebooks are related to the five key thematic areas as follows: (1) Background: Exploring why participants use Ansible and the specific contexts in which they apply it; (2) Advantages: Identifying the key motivations behind adoption; (3) Challenges: Documenting difficulties and limitations; (4) Recommendations: Capturing participants’ wishes; (5) Future: Addressing potential future developments.
For each of these five areas, we developed a codebook. We followed a methodology similar to the one present in prior work (Jackson et al., 2024; Stalnaker et al., 2024; Karlsson et al., 2007), and that is recommended by Human-Centered research guidelines (Lazar et al., 2017). One of the authors developed an initial codebook that captured key aspects of the responses. The initial codebook was shared with a second coder (also an author) , who reviewed and suggested modifications based on a subset of interview data. We repeated this process multiple times until consensus was reached and the codebook was deemed stable.Once we established the final version of the codebook, we double-coded all of the interviews, where one coder systematically applied the codes to all interview responses, and the second coder independently reviewed the applied codes. This review involved cross-checking all coded transcripts to identify inconsistencies or overlooked themes. The second coder made detailed notes on any disagreements. When there was a change to the codebook, for example, a new code, the coding process was restarted. Finally, both coders met to discuss discrepancies and reach a consensus ( 5 and 6 in Figure 3). When attributing quotes, we report the participant ID and the experience of the participant: (B)eginner, (I)ntermediate or (E)xperienced. For example, B1 refers to a beginner participant.
4. RQ1: Common Issues in Ansible Usage
Figure 4. Distribution of posts across topics and their respective themes.
This section presents the challenges encountered by Ansible practitioners, obtained from the topic modeling analysis and interview data. We identified seven primary challenge categories presented in descending order of prevalence.
4.1. Control/Data Flow and Execution
According to our topic modeling analysis, Control/Data Flow and Execution is the most significant theme that represents the challenges Ansible practitioners face. Our approach identified 8,793 posts related to Control Flow and Execution, which represents around 37.9% of all recent posts.
4.1.1. Topic Analysis
Within the Control/Data Flow and Execution theme, Loops represent the most significant challenge for Ansible practitioners (16.4%). As one practitioner stated , “I don’t understand why loops are so hard to do in this language.”. Loops are followed by SSH Connection Issues to hosts (14.9%), Variable Management (14.4%), Conditional Execution (12.4%), and Undefined Variables (12.3%).
Topic Evolution
[…]
4.1.2. Interviews
Participants described various control flow issues. For instance, B1 mentioned frustration with sequential execution slowing runs and a lack of synchronization primitives (n = 2). In more advanced setups, complexities grew around managing dependencies (n = 3).Beyond these granular concerns, a handful of participants reflected on code structure/organization difficulties (n = 5). In contrast, others cited Ansible’s inherently stateless (n = 5) nature as a double-edged sword — simpler to run but lacking built-in features for preserving or comparing states. The push model/agentless (n = 5) approach was both praised and critiqued: praised for minimal overhead yet lamented for complicating continuous enforcement.
4.2. Reliability
Reliability emerged as the second most discussed theme, accounting for approximately 32.3% of all recent posts.
4.2.1. Topic Analysis
Although Error Handling is the most represented topic within the Reliability theme (16.4%), TopicGPT’s interpretation of this topic is quite generic. As a result, despite its description being related to the handling of errors in Ansible playbooks, the topic was frequently used in cases where arbitrary errors occurred. Following Error Handling, the most discussed issues are related to file management (11.5%), permissions (8.1%), JMESPath (5.9%) and the use of shell commands (5.4%).
Topic Evolution
[…]
4.2.2. Interviews
Reliability issues emerged in discussions around debugging. Multiple participants reported that Ansible is hard to debug (n = 10) when playbooks fail mid-run. B2 described moments when “it would get stuck for 15 minutes in a supposedly 1 min script before I forced it to close without any visible changes on the terminal.”. Some lamented the lack of infrastructure debugging tools (n = 5) built into Ansible, while others pointed to bad error messages when modules returned cryptic outputs. As B4 put it, “I actually believe that there’s not a proper way to debugging this [Infrastructure as Code] type of scripts and tools.”.
4.3. Lack of Support
Lack of Support is the second most significant theme, accounting for 32.1% of the recent posts with a total of 7,443 posts.
4.3.1. Topic Analysis
Within this theme, AWX-related issues dominated discussions (20.9%), followed by module execution failures (12.7%), integration issues (10.8%), module usage problems (9.8%), and Windows-related challenges (9.1%). These findings suggest that practitioners encounter persistent challenges with the use of AWX, an open-source project serving as the upstream for Ansible’s enterprise version, along with issues related to module reliability, understanding how to use modules, integrating Ansible playbooks with external systems, and managing Windows hosts.
Topic Evolution
[…]
4.3.2. Interviews
The theme lack of support touched on problems such as missing features that require modules (n = 3), limited OS support (n = 4), and Python version constraints (n = 1). Other participants criticized the lack of support for certain modules (n = 6). As E4 put it “There’s a lack of support for some modules (…). You often have to write your own or hope someone did it.”.
4.4. Usability
Usability represents approximately 16.6% of recent posts, totaling 3,846 posts.
4.4.1. Topic Analysis
The issues related to Jinja2 are by far the most represented, with around 33.1% of the posts within the Usability theme. For instance, in one of the posts, an Ansible practitioner mentioned: “Unfortunately, Jinja2 has been incredibly challenging to work with.”. Issues related to quoting (22.4%), playbook inputs (20.4%), YAML syntax (10.8%), and regex (8.2%) are also prominently discussed. When talking about their issues with YAML, a practioner mentioned: “I find YAML or Ansible’s interpretation of Yaml very frustrating and confusing.”.
Topic Evolution
[…]
4.4.2. Interviews
Ansible’s usability emerged as a significant issue for many participants. It includes user onboarding issues and complaints about complex syntax. Participants expressed that it can be challenging for beginners (n = 10) to grasp the tool’s declarative nature, in particular for those that lacked a coding background (n = 2). Others felt they had to make a mental shift (n = 2), especially if they were used to more imperative scripting, and sometimes it can be hard to pick the right approach (n = 4). Participants also complained about complex syntax (n = 10). Ansible’s reliance on YAML and Jinja introduced the problems of YAML and Jinja complexity (n = 5 and 6). As E5 said, “I once spent 8 hours of work figuring out the escaping on a line of SQL because YAML is so bad at figuring out quotes (…) The hell that is escaping in YAML will never die.”.
4.5. Performance and Scalability
Performance and Scalability represents approximately 7.5% of recent posts, with 1,738 posts discussing these issues.
4.5.1. Topic Analysis
Issues related to dynamic inventories (31.0%) and inventory management (29.8%) are the most frequently discussed topics within Performance and Scalability, followed by performance (17.7%), remote execution (8.5%) and delegation (7.7%).
Topic Evolution
[…]
4.5.2. Interviews
Scalability limitations (n = 8) were frequently mentioned by participants managing large infrastructures. B1 stated: “When you run a playbook for like 100 hosts, it gets really confusing trying to make sense of the messages.”. Participants specifically described problems related to managing inventories (n = 1) and code awareness (n = 4) in large projects. A related source of concern was slow execution (n = 7). While some liked the simplicity of SSH-based workflows, a few individuals cited agentless (n = 4) as detrimental in high-volume scenarios. E4 was particularly direct: “It does not scale when you start to manage 100th of servers, it becomes slow because there is no agent.”.
4.6. Maintenance
Maintenance challenges appear in 683 posts, which corresponds to about 3% of the total of recent posts.
4.6.1. Topic Analysis
Challenges related to deprecated features and settings in Ansible (34.4%), dependency management (23.3%), compatibility issues (19.6%) — which include, for example, supporting legacy operating systems —, different Ansible versions (14.2%) or version controlling Ansible projects (12.3%) are the most relevant topics in the Maintenance theme.
4.6.2. Interviews
Participants noted that certain aspects of Ansible maintenance remain nontrivial. In particular, several mentioned that removing configurations is difficult (n = 3) — as modifying the playbook alone does not guarantee that older changes get undone without explicit tasks. E2 explained that “Doing things like removing config from hosts tends not to work so well in Ansible,”. Configuration drift (n = 5) was also mentioned. As I3 described “Deletes are hard: one cannot merely delete parts of the Ansible playbook and expect those changes to be undone in hosts. Instead, we must change or even add tasks undoing the configuration we want gone.”. Other participants mentioned that Ansible can be hard to maintain (n = 4) and were concerned about deprecations (n = 4).
4.7. Information Gaps
The Information Gaps theme is the least prevalent, accounting for only 465 posts, or about 2% of all recent posts. Of these 465 posts, 464 address the challenges practitioners face with documentation.
4.7.1. Interviews
Another concern revolved around information gaps, particularly documentation that is either overwhelming or incomplete. E2 talked about misleading information: “Another issue is incorrect or misleading answers on forums and newsgroups. Beginners may struggle to distinguish good advice from bad.”. The theme of documentation gaps (n = 8) appeared in several responses, with participants describing the difficulty of locating relevant or clear information among Ansible’s extensive (but sometimes inconsistent) references. Although multiple participants praised Ansible’s core documentation (described in Section 5), they also identified a steep learning curve for interpreting specific advanced.
RQ1: We identified seven core challenge categories — Control/Data Flow and Execution, Reliability, Lack of Support, Usability, Performance and Scalability, Maintenance and Information Gaps — with Q&A emphasizing high-frequency technical issues (e.g., loops, SSH connectivity) and interviews highlighting more nuanced concerns such as stale configurations, misleading community advice, and complex setup processes.
5. RQ2: Factors Influencing Ansible Adoption
This section analyzes how interview participants first encountered Ansible and their reasons for selecting it over alternative IaC solutions. We go over the thematic areas Background and Advantages described in Section 3.3.3.
Initial Exposure to Ansible. Participants described how they discovered Ansible. The most common way was in professional settings (n = 5), where it was already an established tool. Social media platforms also played a notable role in exposure (n = 4), particularly X. As E5 recounted: “In 2016, [someone] posted a tweet about a new configuration management tool (…) I had a look.” Some participants also independently sought out Ansible (n = 3).
Accessibility & Usability. A significant theme in our codebook is the role of Ansible’s accessibility and usability in driving adoption. Several participants described it as Easy to use (n = 12), Easy to onboard to new people (n = 5), and with a Low Barrier to Entry for Setup (n = 6) highlighting how they only needed to master a limited set of concepts—such as basic playbook structures and inventory configuration—before managing a wide range of environments. B1 remarked, “To get things done, I only really needed to understand a small set of concepts regarding how Ansible works (…) ”. Participants also cited Ansible’s Readable syntax (n = 5) and Good Logs (n = 1). E6 highlighted that “[we] were looking into Ansible because it has a nice YAML syntax”.
Tool Qualities and Ecossystem. Beyond accessibility, Ansible was praised for its tool qualities. Participants highlighted its Efficient use of resources, with one participant citing Lightweight resource consumption and three mentioning Scalability. Efficiency was also linked to Cost-Effectiveness, with B4 saying, “Ansible is like a [inexpensive] version [of Kubernetes] for scalability”. Another advantage of Ansible is its Flexibility, since it can be a One-Tool Solution (n = 2) capable of managing heterogeneous environments (e.g., Windows, Linux, or network devices). As E3 puts it, the “One tool that does it all”. Several participants emphasized that Ansible’s Compatibility (n = 4) with various operational requirements and existing workflows set it apart from other tools, and its Agentless (n = 9) model that depends solely on SSH. Idempotence (n = 2) was also considered important for participants, ensuring the same playbook can be applied repeatedly to converge infrastructure into a consistent state. As I3 described: “Some commands are dangerous to invoke twice.” From an Ecosystem standpoint, respondents highlighted features such as Easy to extend (n = 3), Good Tooling (n = 1), and the fact that It’s implemented in Python (n = 2).
Participants frequently framed their rationale in comparative terms, specifically highlighting Ansible’s advantages over alternatives like Puppet, Chef, and Salt. A common comparative advantage cited was the elimination of agent management overhead. E5 provided an illustrative anecdote highlighting Ansible’s comparative operational model advantages: “(…) a place I was working was using Puppet, and the admin was woken up at 3 am from a page, fixed the issue, went to bed, and then Puppet restored the broken configuration, so he woke up again, fixed the Puppet, and then went to bed. Then, was woken up again because Puppet updated itself from git.”.
Community. A third central theme centers on Ansible’s Community. The tool’s Popularity (n = 4) and Community (n = 3) are significant advantages. E1 noted: “Ansible’s advantages (…) are it being a well-supported IaC tool. Since there is a large community around it, it is easier to find solutions to specific problems online.” I1 summarized it as: “Ansible was just the hot topic at the time.” Participants also mentioned that Ansible’s strong market presence often led their organizations or academic institutions to mandate its use (n = 7) (“it was standard by the time I joined the team”, I3). Moreover, if it was not mandatory, it was often Recommended (n = 2) by colleagues or professors.
RQ2: Ansible’s adoption is influenced by its ease of use, low barrier to entry, and agentless design. Participants highlighted how minimal setup requirements and an approachable syntax allowed them to manage diverse infrastructures quickly. They also praised Ansible’s extensibility—particularly its Python-based ecosystem. A strong community presence and widespread popularity were seen as crucial.
6. RQ3: Improving Ansible
This section analyzes what our interviewees would like to see in Ansible and how they see its future. We go over the thematic areas Future and Recommendations described in Section 3.3.3.
6.1. Future of Ansible
When asked about Ansible’s future trajectory, the participants provided various assessments reflecting optimism and caution. A small group (n = 2) expressed expressing concerns that Ansible might become less relevant over time “As more compute ends up in serverless and ephemeral hosts on public and private cloud, Ansible is becoming less relevant.” (E2). The majority of participants took a more Neutral stance, will likely remain a practical choice for specific use cases (n = 5) — as I3 put it “My expectation is that future usage of Ansible will tend towards maintaining legacy systems, and use as a configuration-specific scripting language ”. Others see Ansible retaining its current role (n = 4) in the ecosystem. Other respondents held a Positive outlook, Hoping that the tool will expand in scope and capabilities (n = 2) like described by E5, “I honestly hope that everyone uses. I am teaching my 5-year-old to program, and we’re talking about using Ansible to update his streaming service (mirroring RSS feeds from YouTube) and updating his FDroid repository! I hope that people get empowered to build their systems .”. Others noted that Demand is growing (n = 3) among organizations and peers alike. E2 summarizes their answer by saying “I think the general gist of my opinion is: Ansible is another ‘Worse Is Better’ system. That’s both great and annoying in equal measure, and is why Ansible is as popular as it is.” [^1] This quote inspired the title of our paper..
6.2. Suggestions and Improvements
While participants generally acknowledged Ansible’s strengths, they also offered concrete recommendations to improve.
Documentation and Learning Resources. A recurring concern expressed by participants (n = 8) was the need to improve documentation. E4 described the current official documentation as scattered, noting that it “is poor considering the other configuration tools (take a look at Puppet’s). The websites are not appealing and not properly organized. Strong concepts are scattered over different pages.”. I3 identified specific documentation challenges regarding module organization: “Documentation needs a big revamp: it’s not clear what modules are built-in or part of Ansible Galaxy.”. The participants also called for better onboarding resources (n = 5) and more education (n = 7). As E3 put it “ When you educate beginners, they can identify incorrect information and avoid it.”.
Debugging and Troubleshooting. Several participants highlighted limitations in Ansible’s current debugging capabilities. Better debugging tools (n = 5) and improved error messages (n = 6) were frequently mentioned. E2 stressed that “error messages when plugins fail could be improved,” as they can sometimes be “misleading”. I2 reinforced this point, saying that “more specific error messages would be very helpful. Understanding exactly where an issue presents is the most important part of the debugging process.”. Another concept frequently raised was the lack of rollback-on-failure support (n = 4).
Performance and Scalability. Participant also critiqued performance and mentioned the need to improve speed (n = 6) and improve parallelization (n = 4). As E2 put “Slow runs is an eternal pain point.”. E1 advocated for integrating solutions akin to Mitogen (Mitogen Contributors, 2025), to help optimize remote execution over SSH to reduce overhead and B1 proposed a tag-based approach: “Having the ability to set group tags for tasks, and have groups of tasks be able to run in parallel would be a decent enough solution (tasks with tag A can run in parallel with tasks with tag B).”.
Advanced Programming Capabilities. A group of participants called for the expansion of Ansible’s programming model. While some participants advocated for more high-level language concepts (n = 3) — including real loops, complex conditionals, and synchronization primitives — others (n = 4) stressed state management. E7, for instance, argued for “functionality in Ansible to handle state management effectively” and provided an extensive analysis of Ansible’s state management limitations, saying, “even though Ansible claims to be completely stateless, which is supposed to be its strength, I’m not so convinced. Without state management, it’s challenging to handle certain problems.”.
Participant Experience Levels. While some recommendations, such as providing more education, were popular across experience levels, other priorities varied according to participants’ familiarity with Ansible. Beginners often highlighted the need for better onboarding (n = 2) and rollback on failure (n = 2). In contrast, Experienced users (n = 4) tended to focus on improving speed and other performance enhancements, reflecting the operational demands of managing more extensive infrastructures.
RQ3: Interviewees emphasized the need for better documentation, improved debugging, and enhanced error reporting. Many highlighted performance optimizations, including faster execution and parallelization, while others advocated for rollback support, stronger third-party integrations, and a more user-friendly GUI to streamline workflows.
7. Discussion
This section provides a set of actionable insights and discusses better IaC implications of our work.
7.1. Actionable Insights
This section proposes concrete actionable recommendations for improving the Ansible ecosystem. Each subsection addresses a key domain and references the findings from topic modeling and interviews to explain why these measures are timely and impactful.
7.1.1. Better Debugging Tools
In our analysis of user’s Q&A data, reliability challenges — particularly “Error Handling” — were among the most frequently cited frustrations. Likewise, 10 interview participants described how diagnosing failures becomes particularly difficult when a playbook “gets stuck” with little to no logging feedback. Multiple interviewees lamented the absence of rollback-on-failure or built-in “step-by-step” debugging capabilities that might elucidate the exact point of failure. Although Ansible’s stateless design appeals to many users, it poses challenges for scenarios requiring continuous enforcement or “undo” functionality. While the Q&A data indicates that only a small subset of posts explicitly mention “dependency management” issues (see Figure 4), interviews reveal that participants struggle to remove outdated configurations in real-world environments.
Debugging support could include a “debug mode” generating structured, real-time logs, flagging potential syntax or connectivity issues earlier in the pipeline, and offering recommended fixes. To reconcile stateless simplicity with advanced rollback needs, we propose a configurable “soft state” mode that records minimal metadata—enough to revert undesired changes or compare current system states to a known baseline without imposing excessive overhead. Most debugging currently requires manual intervention, making troubleshooting time-consuming and error-prone. We propose the development of automated repair tools capable of identifying common misconfigurations, detecting inconsistencies in playbooks, and suggesting or even applying corrective actions automatically. Repair mechanisms exist in traditional software development, but their adoption in IaC environments remains limited (Weiss et al., 2017).
7.1.2. Higher-Level Language Concepts
Although Ansible’s declarative approach and reliance on YAML/Jinja2 can simplify basic configuration tasks, our analysis of Q&A data found syntax issues with Jinja and other parts of Ansible were the main challenges of the Usability theme. Several interviewees (n = 10), mentioned the limitations in expressing complex logic. They described frustrations with YAML syntax quirks (especially for loops and conditionals), cumbersome variable handling, and a lack of “proper” language features for advanced use cases.
Addressing these challenges involves introducing higher-level language concepts. One possible approach is to bridge YAML with Python-based enhancements, allowing better support for loops, conditionals, or data processing without breaking Ansible’s declarative philosophy. This aligns with ongoing developments in Ansible’s SDK, which provides APIs for extending Ansible’s functionality using Python. Additionally, solutions like Pulumi (Pulumi, 2025) present an alternative paradigm where IaC is directly implemented in imperative programming languages rather than YAML-like configurations. While Ansible remains firmly declarative, allowing for optional hybrid workflows where advanced users can script automation logic in Python would reduce the need for error-prone YAML workarounds. We argue that the goal is not to replace YAML entirely but to augment it for scenarios that exceed basic configuration.
7.1.3. Improve Performance
Our topic modeling analysis revealed a substantial volume of posts on slow execution (17.7% of the posts related to Performance and Scalability). In interviews, seven participants reported frustrations with performance bottlenecks, particularly when managing hundreds of hosts. They cited sequential task execution, agentless overhead, and limited parallelization features as key obstacles slowing down their workflows.
Participants and our Q&A data analysis converge on the need for better parallelization and reduced overhead in large deployments. An alternative tool is Jetporch (Jetporch Contributors, 2025), a Rust-based Ansible alternative that aims to address performance limitations by avoiding Python (and its inherent execution overhead) and optimizing execution. Jetporch’s design ideas show key areas where Ansible could improve. Ansible should undergo refactoring to introduce parallelization improvements, optimize SSH connect, and enable partial execution in a more efficient manner.
7.1.4. Improve documentation
Documentation problems were mentioned more often in the interviews than in our Q&A data (only 2% of the recent posts). In our survey, eight participants flagged documentation gaps as a pain point. Users commonly struggle to locate authoritative references for advanced features or specialized modules. Interviewees recounted sifting through outdated or misleading community-contributed tutorials.
Good documentation is one of the things that our participants liked about Ansible, so we argue that Ansible should capitalize on this fact by iterating over the current documentation, centralizing official resources, clarifying built-in vs. Galaxy modules, and providing version-specific references. Beyond traditional methods of documentation overhaul, generative AI tools could dynamically address knowledge gaps. For instance, LLMs could identify deprecated or outdated sections by continuously parsing official documentation and user-contributed content, and flagging inaccurate sections, automatically suggesting updates. Similarly, a tool like this could produce playbook snippets or configuration examples for less common scenarios on demand.
7.2. Future Work and Broad IaC Implications
Our findings highlight key challenges and opportunities for Ansible’s evolution and offer broader insights into the trade-offs inherent in IaC tool design.
Mixed-Methods Framework.
We observe differences between publicly discussed topics and the more subjective issues revealed through direct interviews. While thematic areas such as “Control Flow” or “Lack of Support” dominate forum posts, other challenges — like documentation gaps or configuration drift — appear more in interviews. For instance, information gaps and misleading advice may not prompt a direct question on a forum. Practitioners may not even realize they are facing an information gap, nor might they post about it explicitly. Instead, they continue troubleshooting, or they might consult colleagues offline. In interviews, however, participants openly discussed these intangible pain points — highlighting subjective experiences. We argue that qualitative insights should be used with quantitative or text-mining approaches to provide a holistic understanding of IaC adoption barriers. We provide a straightforward and replicable methodology that can be used for other major IaC platforms (e.g., Chef, Puppet) to yield more holistic, data-driven evaluations of challenges that practitioners face.
Balancing Usability and Scalability.
User experience seems to vary by expertise level. Beginners emphasize documentation clarity, onboarding simplicity, and accessible debugging tools. Meanwhile, advanced practitioners advocate for performance optimizations (e.g., concurrency) and advanced programming capabilities. This suggests that IaC tools should provide tiered user experiences — offering simplified workflows for beginners while exposing advanced performance optimizations and parallelization capabilities for experts. Striking this balance seems to be important to ensure long-term adoption across skill levels.
Documentation Gaps are a Universal Problem.
Finally, documentation gaps and information drift are persistent themes across IaC platforms, even in Ansible. While well-developed official documentation can lower the barrier to entry (Robillard, 2009; Aghajani et al., 2019), many tools struggle to keep pace with rapid feature expansion and a growing user base (Aghajani et al., 2020). Establishing centralized, consistently updated repositories of best practices and official modules is important for sustained usability.
8. Threats to Validity
Threats to Internal Validity.
A threat to internal validity arises from the use of LLMs through TopicGPT. LLMs are prone to hallucination and can generate unreliable results, potentially affecting the precision of our findings. To mitigate this threat, we perform manual validation in two stages: (1) during the generation of the topic list, and (2) during the assignment of topics to individual posts. Also, conducting semi-structured interviews introduces risks such as interviewer bias, where the phrasing or emphasis of questions might influence participant responses. To minimize this threat, we followed a consistent interview protocol and performed pilots before deploying the study.
Threats to External Validity.
A threat to external validity is that practitioners can discuss Ansible issues on alternative Q&A platforms that we did not analyze, potentially limiting the representativeness of our findings. To mitigate this, we collected posts from three distinct platforms. Another threat is that while some of our findings may generalize to other IaC tools, particularly configuration management tools, more research is needed to confirm their applicability. Future studies should explore other IaC tools to assess the generalizability of our results. Regarding the interviews, we recruited participants primarily through online channels, so our sample may not be representative of the broader Ansible user community. Self-selection bias may have resulted in participants who are more engaged with the tool or have stronger opinions about it. Furthermore, our sample’s distribution across expertise levels may not accurately reflect the actual distribution in the community.
9. Related Work
To identify challenges that practitioners face in IaC, Rahman et al. analyzed 2,758 Puppet-related questions from Stack Overflow (Rahman et al., 2018). Through qualitative analysis, the authors identified 16 major categories of questions. Many of these categories, including Syntax Error and Filesystem, are also observed in our study. While Rahman et al.’s study focuses on Puppet, our study focuses on Ansible, which is as of 2024 a more widely used technology for configuration management of services. Furthermore, Rahman et al. analyzed questions posted between January 2010 and December 2016, while our interviews and posts reflect the scenario as of 2024.
Guerriero et al. conducted 44 semi-structured interviews with IaC practitioners to understand how practitioners adopt and develop IaC, the tools available within the IaC ecosystem along with their advantages and limitations, and the specific needs and challenges that practitioners face when working with IaC (Guerriero et al., 2019). Similarly to our study, they highlight issues such as difficulties with testing and debugging. However, we provide a more granular and in-depth analysis, as well as the identification of unique issues of Ansible, such as issues with loops and Jinja.
Begoug et al. collected 52,692 questions and 64,078 responses from Stack Overflow to investigate challenges developers encounter with IaC (Begoug et al., 2023). The dataset includes questions related to Ansible, Pulumi, Terraform, CloudFormation, and other IaC technologies. The authors grouped questions into specific topics using the Latent Dirichlet Allocation (LDA) method optimized with a Genetic Algorithm (GA) for parameter fine-tuning. Our analysis of the questions of practitioners follows a similar approach. However, we collected data from two additional sources: Reddit and the Ansible forum. Additionally, instead of using LDA, we use TopicGPT for topic modeling, which not only provides interpretable topics, but also produces topics that align better with human categorizations than LDA (Pham et al., 2024). Finally, although we both found common topics such as File Management and Templating, due to our focus on Ansible, we found more granular and specific challenges faced by practitioners such as Jinja and AWX issues.
To the best of our knowledge, the other only mixed methods study related to IaC challenges is the one conducted by Tanzil et al. (Tanzil et al., 2023). However, their study focuses on a broader scope, examining DevOps challenges in general, with IaC being one of the categories included. They used LDA to perform topic modeling on approximately 174,000 Stack Overflow posts, identifying 23 distinct topics. Then these topics were grouped into four categories, IaC being one of them, encompassing seven specific topics. Finally, the authors validated and extended their findings with a survey of 21 professional DevOps practitioners. Although there is some overlap in the findings between our study and that of Tanzil et al. (for example, both studies identify challenges related to File Management and Syntax), we focus exclusively on challenges related to Ansible, which provides a much narrower scope with different insights such as the issues on control flow, such as difficulties with loops and conditionals. We also employ a different topic modeling approach and focus on interviews over surveys.
10. Conclusion
This paper combined a large-scale quantitative analysis of 59,157 online Ansible discussions with 16 practitioner interviews to map out the most significant challenges, contextualize their real-world impact, and propose targeted solutions. We found issues related to control flow, performance bottlenecks, limited debugging capabilities, and inconsistent documentation, highlighting the tensions among usability, flexibility, and scalability within Ansible’s ecosystem. We offer concrete recommendations and show how Ansible can better evolve to serve novice users and large-scale deployments. These contributions, alongside broader reflections on community-driven knowledge, provide a replicable roadmap for assessing and improving IaC solutions.
Acknowledgements.
We would like to thank Claúdia Mamede, who helped us design our main diagram. This work was supported by FCT, Fundação para a Ciência e a Tecnologia under grants PRT/BD/153739/2021 and BD/04736/2023, and projects UIDB/50021/2020 (DOI: 10.54499/UIDB/
50021/2020), LA/P/0063/2020 (DOI: 10.54499/LA/P/0063/2020), and 2024.07411.IACDC.